### v1

# Optimal performance of a three stroke heat engine in the microscopic regime

Abstract
In the microscopic regime, we consider a stroke-based heat engine that utilizes the temperature difference between two heat baths to produce work using a working body. The working body of the engine aims to extract heat from the hot heat bath, generate work, and discharge the surplus heat into the cold heat bath through the successive execution of three strokes. We discuss the thermodynamic framework for such an engine that involves any arbitrary-dimensional working body. Next, our investigation focuses on the smallest version of this engine, which involves a two-level working body, resulting in a closed expression for optimal work production and efficiency. Furthermore, we investigate the possibility of realizing such an engine within the Jaynes-Cummings model, gaining insights into optimal work production and efficiency. Finally, we scale down the engine, taking into account finite-dimensional heat baths, and obtain the expressions for optimal work and efficiency in this scenario.
I Introduction
Within the domain of thermodynamics, heat engines play a pivotal role by converting thermal energy into work. The study of heat engines not only propels technological advancements in energy conversion but also deepens our understanding of nature’s fundamental physical laws. However, the advent of quantum mechanics has ushered in a new era of insight, particularly concerning the foundational underpinnings of thermodynamics on a microscopic scale. This captivating avenue of inquiry traces its origin back to the 1950s, when investigations into the thermodynamic properties of lasers commenced [1, 2, 3]. Since then, a resounding interest has flourished in comprehending the mechanisms of quantum thermal engines as well as quantum thermal machines in general [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], and in formulating microscopic thermodynamic frameworks to characterize their operations [22, 23, 24, 25, 26, 27]. Recently, researchers have realized the microscopic thermal heat engines in the lab using a trapped single calcium ion as a working body [28, 29, 30, 31, 32], as well as in superconducting circuits [33], nitrogen vacancy centers in diamonds [34] atomic [35] and phononic systems [36], as well as single-electron transistors [37, 38].
Microscopic heat engines can be described as thermodynamic devices that involve small-sized components. In particular, the microscopic heat engines described in the literature involve working bodies that are finite-dimensional and can be as simple as two-level systems [23, 15]. Furthermore, the analysis of the microscopic heat engines has been extended further by considering heat baths of finite size [39, 40, 41, 42, 43]. Based on their way of functioning, the heat engines considered in the microscopic regime can be broadly classified into two classes: continuous engines [11, 7, 44, 45, 46, 47] and stroke-based engines [48, 49, 50, 51]. In a continuous engine, the working body is always in contact with the hot and cold heat baths. As a consequence, the engine withdraws heat from the hot heat bath and converts it into work in a continuous and simultaneous manner. On the other hand, in a stroke-based heat engine, the working body exchanges heat and converts it into work in discrete strokes that are well separated in time. Therefore, in stroke-based engines, the working body repeatedly couples between the hot and cold baths using subsequent strokes.
The complete characterization of a continuous heat engine necessitates the consideration of various models of heat baths and couplings. These models are typically formalized through master equations, which, in turn, depend on multiple parameters such as coupling constants, bath spectral densities, etc [52, 53]. Consequently, optimizing the work extraction and efficiency of a continuous heat engine proves to be a formidable task. Despite the inherent complexity, a more accessible understanding of the basic operations of these continuous heat engines can be attained by considering stroke-based heat engines. Importantly, this approach doesn’t require detailed knowledge of the dynamics. By embracing this perspective, one can acquire significant knowledge about the fundamental limitations, optimal work, and efficiency of continuous thermal machines.
In light of this, in this paper we examine a stroke-based heat engine that operates in the microscopic regime. By microscopic, we are referring to the fact that only two-level systems have been taken into account as working bodies. In the first stroke the working body aims to withdraw heat from the hot heat bath, followed by the second and third stroke where the working body converts the heat into the work and disposes the extra amount of heat into cold heat bath and completes the cycle. In order to prevent any unaccounted energy sources from contributing to the work done by the engine, we assume that the interaction between the working body and heat baths is energy-conserving. In other words, the engine generates work by harnessing the temperature difference between the two heat baths. This led us to describe the transformation of the state of the working body via a thermal operation [54, 26, 55]. The process of extracting work is represented by a unitary process, in which the working body of the engine converts a portion of the heat extracted from the hot heat bath into work, while the remaining heat is deposited to a cool heat bath. We provide the thermodynamic framework for this specific heat engine. In particular, we show that the efficiency of such an engine is always upper bounded by the Carnot efficiency. In this work, we introduce a straightforward way to optimize the efficiency of such an engine in a more generalized setting, where the range of thermal operations with respect to a heat bath is restricted. Considering the restricted range of the thermal operations, we can investigate the feasibility of implementing this engine in a more realistic scenario, such as the Jaynes-Cummings interaction. Furthermore, we extend our study to a scenario in which the heat bath’s dimension is finite.
The paper is organized as follows. In Sec. II, we describe the construction of the stroke-based heat engine. We outline the individual strokes that constitute the thermodynamic cycle for this specific three-stroke heat engine. The mathematical framework that describes the characteristics of the strokes is also presented. Subsequently, we establish the thermodynamic framework to precisely describe the process of extracting work and exchanging heat in this stroke-based heat engine. In Sec. III, we calculate the maximum work production and efficiency. The results are obtained by considering a general range of thermal operations that describe the transformation of the working body when it comes into contact with the heat bath. Using the general expressions of optimal work production and efficiency, we examine the work production and efficiency for specific instances, such as when the working body’s interaction with heat baths is modelled by the Jaynes-Cummings interaction and the heat bath’s dimension is finite (modelled by a truncated harmonic oscillator). Finally, we conclude in Sec. IV.
II Description of a stroke-based discrete engine
II.1 Construction of the engine
Any stroke-based heat engine operates in a sequence of strokes that allows it to consume heat from a hot heat bath and convert it into work in subsequent strokes. The second law of thermodynamics asserts that it is impossible to convert all the heat that has been drawn from the hot heat bath to work. Thus, in order to work in a cyclic manner, the engine needs to release a fraction of the consumed heat into the cold heat bath after work extraction in the final stroke. Here, we consider a model of a stroke-based engine in the microscopic regime that requires three subsequent strokes to complete the thermodynamic cycle. The heat engine consists of the following key components:
  1. 1.
    A hot heat bath is thermalized at an inverse temperature $\beta_{H}$, which serves as a source of heat that requires to produce work. We label this component by $H$.
  2. 2.
    A cold heat bath is thermalized at inverse temperature $\beta_{C}$ and works as a heat sink for dumping the remainder of the consumed heat from the hot heat bath after work has been extracted. We denote this component by $C$.
  3. 3.
    A working substance or working body represented by a $d$-level quantum system that controls the flow of energy among other components of the engine. We label this component by $S$.
Therefore, the total free Hamiltonian of the whole system is
$H_{\text{free}}=H_{S}+H_{H}+H_{C},$ (1)
where $H_{S}$, $H_{H}$, and $H_{C}$ denote the Hamiltonian of the working body, hot heat bath, and cold heat bath, respectively. In the start, we assume all the components of the engine are uncorrelated, which allows us to write the full initial state $\rho_{\text{in}}$ of the engine as
$\rho_{\text{in}}=\rho_{S}\otimes\gamma_{H}\otimes\gamma_{C},$ (2)
where $\gamma_{H}$ and $\gamma_{C}$ denote the thermal states of the hot and cold heat baths, i.e.,
$\gamma_{H}=\frac{e^{-\beta_{H}H_{H}}}{\text{Tr}(e^{-\beta_{H}H_{H}})};\quad\gamma_{C}=\frac{e^{-\beta_{C}H_{C}}}{\text{Tr}(e^{-\beta_{C}H_{C}})}.$ (3)
Whereas, $\rho_{S}$ represent the initial state of the working body. In the following section, we shall discuss the functioning of the engine.
II.2 Characterization of the strokes of the engine
This section focuses on the various strokes that allow the engine to operate. The engine operates on three major strokes: i) heat stroke, ii) work stroke, and iii) cold stroke. While the engine involves energy-conserving stroke operations, this does not result in a fully autonomous engine since it needs external control to carry out the various steps. Nevertheless, the engine is an advancement towards creating an autonomous machine as it preserves total energy. To put it simply, the process of enabling or disabling interactions within the system does not impact the flow of energy into or out of the system. Following this, we describe the workings of the engine in more detail.
II.2.1 Heat stroke
This is the initial stroke of the engine, in which the working body interacts with the hot heat bath at thermal equilibrium using energy-saving unitary $U_{S,H}$, i.e, $[U_{S,H},H_{S}+H_{H}]=0$. Thus, the transformation of the state $\rho_{S}$ is governed by a thermal operation ${\cal E}_{H}$ [56], i.e.,
$\rho_{S}\rightarrow{\cal E}_{H}(\rho_{S}):=\text{Tr}_{H}\big{(}U_{S,H}(\rho_{S}\otimes\gamma_{H})U_{S,H}^{\dagger}\big{)}.$ (4)
The aim of this stroke is to consume heat from the hot heat bath, which increases average energy and induces non-passivity in the working body. In accordance with classical thermodynamics, where heat is defined as the amount of energy transferred to the system from the hot heat bath, we define the amount of heat transferred, denoted as $Q_{H}$, as the change in the system’s average energy caused by its interaction with the hot heat bath, which can be expressed as:
$Q_{H}=\text{Tr}\Big{(}\big{(}H_{S}\big{(}\mathcal{E}_{H}(\rho_{S})-\rho_{S}\big{)}\Big{)}.$ (5)
Here, $\mathcal{E}_{H}(\rho_{S})$ represents the final state of the working body after the interaction with the hot heat bath.
II.2.2 Work stroke
In the second stroke of the engine, the working body, which resulted from the first stroke, attempts to produce work using the heat it drew from the first stroke. The engine extracts work using a unitary operation that transforms the working body’s state from the first stroke to one with lower average energy. Therefore, this stroke transforms the working body’s state as follows:
$\mathcal{E}_{H}(\rho_{S})\rightarrow V(\mathcal{E}_{H}(\rho_{S}))V^{\dagger}.$ (6)
The amount of extracted work is given by
$W=\text{Tr}\Big{(}H_{S}\big{(}\mathcal{E}_{H}(\rho_{S})-V(\mathcal{E}_{H}(\rho_{S}))V^{\dagger}\big{)}\Big{)},$ (7)
where $V$ is any arbitrary unitary. The above expression clearly shows that the unitary, which orders the eigenvalues of the density matrix $\mathcal{E}_{H}(\rho_{S})$ in decreasing order, will maximize the extracted amount. In other words, the expression of work done by the engine as given in Eq. (7) will be maximized for unitary, which leads to the passive version of the state $\mathcal{E}_{H}(\rho_{S})$ [8]. This maximum amount of the extracted work is defined as ergotropy of the state $\rho_{S}$, i.e.,
$W$ $=$ $\text{Tr}\Big{(}H_{S}\big{(}\mathcal{E}_{H}(\rho_{S})-V(\mathcal{E}_{H}(\rho_{S}))V^{\dagger}\big{)}\Big{)}$ (8)
$\leq$ $\max_{V\in\mathbb{U}(d)}\text{Tr}\Big{(}H_{S}\big{(}\mathcal{E}_{H}(\rho_{S})-V(\mathcal{E}_{H}(\rho_{S}))V^{\dagger}\big{)}\Big{)}$ (9)
$=$ $\text{Tr}\Big{(}H_{S}\big{(}\mathcal{E}_{H}(\rho_{S})-\mathcal{E}_{H}^{\downarrow}(\rho_{S})\big{)}\Big{)}=R(\mathcal{E}_{H}(\rho_{S})),$ (10)
where $\mathcal{E}_{H}^{\downarrow}(\rho_{S})$ denotes the passive version of the state $\mathcal{E}_{H}(\rho_{S})$, $R(\mathcal{E}_{H}(\rho_{S}))$ denotes the ergotropy of the state $\mathcal{E}_{H}(\rho_{S})$ and $\mathbb{U}(d)$ is the set of $d$-dimensional unitary [8, 57]. Next, we shall describe the final stroke of the working body.
II.2.3 Closing the thermodynamic cycle
In the final stroke of the engine, the state of the working body that results from the work stroke $V{\cal E}_{H}(\rho_{S})V^{\dagger}$ interacts with the cold heat bath in thermal equilibrium via an energy conserving unitary $U_{S,C}$, i.e., $[U_{S,C},H_{S}+H_{C}]=0$. This transformation corresponds to a thermal operation with respect to the cold heat bath denoted by ${\cal E}_{C}$, i.e.,
$\mathcal{E}_{C}\big{(}V{\cal E}_{H}(\rho_{S})V^{\dagger}\big{)}:=\text{Tr}_{C}\Big{(}U_{S,C}\big{(}(V{\cal E}_{H}(\rho_{S})V^{\dagger})\otimes\gamma_{C}\big{)}U_{S,C}^{\dagger}\Big{)}.$ (11)
The goal of this stroke is to return the working body to its initial state, i.e.,
$\mathcal{E}_{C}\big{(}V{\cal E}_{H}(\rho_{S})V^{\dagger}\big{)}=\rho_{S},$ (12)
that results in the closing of the thermodynamic cycle. Our aim is to figure out the cyclic state $\rho_{S}$ that satisfies the relation given in Eq. (12).
We define cold heat as the amount of heat that the working body releases into the cold heat bath, expressed as follows:
$Q_{C}=\text{Tr}\Big{(}H_{C}\big{(}\rho_{S}-V{\cal E}_{H}(\rho_{S})V^{\dagger}\big{)}\Big{)}.$ (13)
This step is critical for the entire setup to function as an engine. More generally, this stroke encodes the manifestation of the second law for this heat engine.
II.3 The open cycle engine
The open cycle engine introduced in Ref. [20] represents a subclass of the discussed three-stroke engine. This kind of engine starts with a working body that is thermalized at the inverse temperature of the cold heat bath. Therefore, the final stroke of the engine simply reduces to thermalization at the inverse temperature of the cold bath, which results in the completion of the cycle. Our research aims to bridge the understanding between an open-cycle engine and a generic three-stroke engine that we have introduced previously, particularly when the working body comprises two-level systems. This connection significantly simplifies the optimization of efficiency and work for the three-stroke engines with a two-level working body. The comparison between these two engines has been illustrated in Fig. 1.
Figure 1: The above figure depicts the equivalence between the proposed three-stroke engine and open cycle engine introduced in Sec. II.3. On the left panel we have shown the proposed three stroke engine where the first stroke and the third stroke aiming for withdrawing heat from the hot bath and disposing the excess heat in the cold bath, whereas the production of work happened during the second stroke. The first and third strokes are characterized by the thermal operations at the hot and cold inverse temperatures $\beta_{H}$ and $\beta_{C}$, respectively. This proposed engine is equivalent to an open cycle engine where the third stroke is characterized by thermalization at some effective inverse temperature $\beta^{v}_{C}$ as shown in the right panel. This equivalence is capturing the underlying idea behind the proof of theorem 2.
II.4 Mathematical tools
This section provides a brief overview to the mathematical tools associated with thermal operations that are important to evaluate the performance of the engine. A thermal operation can be defined as a CPTP map ${\cal E}$, which transforms a system’s state $\rho$ as
$\mathcal{E}(\rho)=\text{Tr}_{E}\Big{(}U(\rho\otimes\gamma_{E})U^{\dagger}\Big{)},$ (14)
where $U$ satisfies the relation $[U,H_{S}+H_{E}]=0$ with $H_{S}$, $H_{E}$ are the Hamiltonians of the system and bath, respectively, and $\gamma_{E}=\frac{e^{-\beta H_{E}}}{\text{Tr}(e^{-\beta H_{E}})}$ is the thermal state of the bath that is thermalized at an inverse temperature $\beta$.
The central task is then described by the following interconversion problem: what is the set of final states that can be achieved from a given state $\rho$ via thermal operations? This set is convex and referred to as the thermal cone of the state $\rho$, denoted by ${\cal T}(\rho)$, which follows straightforwardly from the convexity of thermal operation [58]. Characterizing the necessary and sufficient conditions for the interconversion problem in general is challenging and unknown beyond dimension two [59, 60]. Nonetheless, the seminal work by Horodecki and Oppenheim in Ref. [56] characterizes the necessary and sufficient conditions for the state interconversion problem when $[\rho,H_{S}]=0$. These states are referred to as energy-incoherent states as they do not possess coherence in the energy eigenbasis. For a given energy incoherent state $\rho$, the thermal cone ${\cal T}(\rho)$ is reduced to a polytope, i.e., the set will have a finite number of extremal points. In this case, we refer the set ${\cal T}(\rho)$ as the thermal polytope. The thermal operation that transforms the initial energy incoherent state $\rho$ to an extremal point of the thermal polytope ${\cal T}(\rho)$ is called an extremal thermal operation. Later, we shall see that these extremal points of the set ${\cal T}(\rho)$ and the extremal thermal operations play a crucial role in optimizing the performance of the three stroke engine.
Let us briefly describe the necessary and sufficient conditions for the interconversion problem for energy-incoherent states. In order to do so, we associate a probability vector $\bm{p}$, called the population vector, with an energy incoherent state $\rho$ such that $p_{i}=\langle E_{i}|\rho|E_{i}\rangle$, where $|E_{i}\rangle$ is an eigenstate of Hamiltonian $H_{S}$ with eigenvalue $E_{i}$. Then, an energy incoherent state $\rho$ can be converted to another energy incoherent state $\sigma$ via a thermal operation ${\cal E}$ iff there exists a stochastic matrix $T^{{\cal E}}$ such that
$T^{{\cal E}}\bm{p}=\bm{q}\quad\quad T^{{\cal E}}\bm{\gamma}=\bm{\gamma},$ (15)
where $\bm{p}$, $\bm{q}$ and $\bm{\gamma}$ are the population vectors associated with $\rho$, $\sigma$, and $\gamma$, respectively. Here, $\gamma$ is the Gibbs state of the system. We call such a stochastic matrix $T^{{\cal E}}$ a Gibbs-stochastic matrix or thermal process [55, 61]. Employing the concept of relative majorization, one can prove that the existence of such a Gibbs-stochastic matrix is equivalent to the thermomajorization condition, which we have described in Appendix B [55, 62, 63]. The set of all thermal processes for a given dimension is a polytope. Among the extremal thermal processes, we shall be interested in those thermal processes that correspond to the extremal thermal operation.
Let us illustrate the above-mentioned concepts for a two-level energy incoherent state, which will be crucial for optimizing the performance of the heat engine. For an energy incoherent initial state
$\rho=\begin{pmatrix}p&0\\ 0&1-p\end{pmatrix},$ (16)
with Hamiltonian
$H_{S}=\omega|1\rangle\!\langle 1|,$ (17)
we associate the population vector
$\bm{p}=\begin{pmatrix}p&1-p\end{pmatrix}^{T}.$ (18)
The extremal thermal process in dimension $2$ is given by [26]:
$\quad\Bigg{\{}\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\begin{pmatrix}1-e^{-\beta\omega}&1\\ e^{-\beta\omega}&0\\ \end{pmatrix}\Bigg{\}}.$ (19)
Therefore, the thermal polytope for the state $\rho$ in Eq. (16) is given by
${\cal T}(\rho)=\text{Conv}\Bigg{\{}\rho,\begin{pmatrix}1-pe^{-\beta\omega}&1\\ pe^{-\beta\omega}&0\end{pmatrix}\Bigg{\}},$ (20)
where $\text{Conv}\{A\}$ denotes the convex hull of the set $A$. Thus, any arbitrary element of the set ${\cal T}(\rho)$ can be expressed as:
$\lambda\begin{pmatrix}1-pe^{-\beta\omega}&1\\ pe^{-\beta\omega}&0\end{pmatrix}+(1-\lambda)\rho$ (21)
with $\lambda\in[0,1]$.
Here, we would like to make an important remark. When we define thermal operations, we suppose that our system of interest is in touch with a very large bath that has an effectively infinite volume. In such an ideal scenario, all possible states within the set ${\cal T}(\rho)$ can be realized. In a more practical context, various restricted subsets of states emerge within the thermal polytope ${\cal T}(\rho)$ as a result of constraints involving the feasibility of thermal operations. In such cases, the range of $\lambda$ in Eq. (21) is more generally, confined between
$[0,\lambda^{\text{max}}_{\beta}],\;\text{where}\;0\leq\lambda^{\text{max}}_{\beta}\leq 1.$ (22)
In this paper, two particular subsets that will be of interest to us are as follows:
  1. 1.
    The set of states that can be achieved from a given state of two-level systems through thermal operations in the presence of a finite heat bath, which can be realized with a truncated harmonic oscillator. [64].
  2. 2.
    The set of states which are achievable from a given state of two-level systems via thermal operations that can be realized via rotating wave approximated Jaynes-Cummings interaction [65].
After explaining all of the necessary concepts related to the functioning of the three-stroke engine, we will now proceed by describing the main results of this work.
III Main Results
III.1 Thermodynamic framework
In this section, we shall develop the thermodynamic framework of the proposed heat engine. From the description of the three stroke heat engine that we have discussed in the previous section (see Ref. II), we can summarize the thermodynamic cycles of the three stroke engine as
$\rho_{S}\rightarrow{\cal E}_{H}(\rho_{S})\rightarrow V{\cal E}_{H}(\rho_{S})V^{\dagger}\rightarrow{\cal E}_{C}\big{(}V{\cal E}_{H}(\rho_{S})V^{\dagger}\big{)}=\rho_{S}.$ (23)
From the flowchart of Eq. (23), we can write the total change in average energy as
00 0 $=$ $E(\rho_{S})-E(\rho_{S})$ (24)
$=$ $E(\rho_{S})-E({\cal E}_{H}(\rho_{S}))+E({\cal E}_{H}(\rho_{S}))-E(V{\cal E}_{H}(\rho_{S})V^{\dagger})$
$+E(V{\cal E}_{H}(\rho_{S})V^{\dagger})-E(\rho_{S})$
$=$ $-Q_{H}+W-Q_{C},$
where $E(\cdot)=\text{Tr}(H_{S}\;\cdot)$ denotes average energy, and in the last line, we have used Eqs. (5), (7), and (13). Equation (24) establishes the first law of thermodynamics as the manifestation of the conservation of energy
$W=Q_{H}+Q_{C},$ (25)
which allows us to express the work done by the engine as the net average flow of heat. Then, the efficiency of the engine is defined as
$\eta:=\frac{W}{Q_{H}}=\frac{Q_{C}+Q_{H}}{Q_{H}}=1+\frac{Q_{C}}{Q_{H}}.$ (26)
Next, we demonstrate that the efficiency of the engine is upper bounded by the Carnot efficiency, therefore establishing the second law of thermodynamics. The result is stated in the following theorem.
Theorem 1.
The efficiency of the three-stoke engine is bounded by the Carnot efficiency,
$\eta\leq 1-\frac{\beta_{H}}{\beta_{C}}:=\eta_{\textnormal{Carnot}}.$ (27)
The proof of the theorem can be found in appendix A. This establishes the thermodynamic framework of the three-stroke heat engine. We shall now optimize the work production and efficiency of the three-stroke engine.
III.2 Optimal performance of the three stroke engine
In this section, we characterize the optimal efficiency and work produced by the three-stroke engine.
Theorem 2.
Consider a three-stroke engine with a two-level system as the working body with Hamiltonian $H_{S}=\omega|1\rangle\langle 1|$ and the range of $\lambda$ in Eq. (22) is restricted between 00 to $\lambda^{\textnormal{max}}_{H}$ and $\lambda^{\textnormal{max}}_{C}$ for the first and third strokes, respectively. Then, the optimal work produced by the engine and the optimal efficiency are given by
$p$ $=$ $\frac{1-\lambda^{\textnormal{max}}_{H}(1-\lambda_{C}^{\textnormal{max}})-\lambda_{C}^{\textnormal{max}}(1-\lambda^{\textnormal{max}}_{H})e^{-\beta_{C}\omega}}{2-\lambda_{C}^{\textnormal{max}}(1-\lambda_{H}^{\textnormal{max}})-\lambda_{H}^{\textnormal{max}}-\lambda_{C}^{\textnormal{max}}(1-\lambda_{H}^{\textnormal{max}})e^{-\beta_{C}\omega}-\lambda_{H}^{\textnormal{max}}(1-\lambda_{C}^{\textnormal{max}})e^{-\beta_{H}\omega}+\lambda_{H}^{\textnormal{max}}\lambda_{C}^{\textnormal{max}}e^{-(\beta_{C}+\beta_{H})\omega}},$ (28)
$W_{\textnormal{max}}$ $=$ $1-2\lambda^{\textnormal{max}}_{H}+\frac{2\Big{(}\lambda_{H}^{\textnormal{max}}e^{-\beta_{H}\omega}-(1-\lambda_{H}^{\textnormal{max}})\Big{)}\Big{(}1-(1-\lambda_{C}^{\textnormal{max}})\lambda_{H}^{\textnormal{max}}-\lambda_{C}^{\textnormal{max}}e^{-\beta_{C}\omega}(1-\lambda_{H}^{\textnormal{max}})\Big{)}}{2-\lambda_{C}^{\textnormal{max}}(1-\lambda_{H}^{\textnormal{max}})-\lambda_{H}^{\textnormal{max}}-\lambda_{C}^{\textnormal{max}}e^{-\beta_{C}\omega}(1-\lambda_{H}^{\textnormal{max}})-\lambda_{H}^{\textnormal{max}}e^{-\beta_{H}\omega}(1-\lambda_{C}^{\textnormal{max}})+\lambda_{C}^{\textnormal{max}}\lambda_{H}^{\textnormal{max}}e^{-(\beta_{C}+\beta_{H})\omega}},$
$\eta_{\textnormal{max}}$ $=$ $1-\frac{\lambda_{C}^{\textnormal{max}}\Big{(}1-\lambda_{H}^{\textnormal{max}}e^{-\beta_{H}\omega}-(1-\lambda_{H}^{\textnormal{max}})e^{-\beta_{C}\omega}\Big{)}}{\lambda^{\textnormal{max}}_{H}\Big{(}e^{-\beta_{H}\omega}-(1-\lambda^{\textnormal{max}}_{C})-\lambda_{C}^{\textnormal{max}}e^{-(\beta_{H}+\beta_{C})\omega)}\Big{)}}.$ (30)
The complete proof of this theorem can be found in Appendix C which requires the concept of $\beta$-ordering that we have described in Appendix B [26, 55].
Let us analyze the scenario where one can perform any thermal operation at inverse temperature $\beta_{C}$ and $\beta_{H}$ in the first and third strokes, respectively. This implies $\lambda^{\text{max}}_{H}=1$ and $\lambda^{\text{max}}_{C}=1$. In that scenario, the optimal work produced by the engine and the optimal efficiency from Eqs. (29) and (30) are reduced to [66]
$W_{\text{max}}$ $=$ $\omega\Big{(}\frac{2e^{-\beta_{H}\omega}}{1+e^{-(\beta_{H}+\beta_{C})\omega}}-1\Big{)}\,,\mbox{and}$
$\eta_{\text{max}}$ $=$ $1-\frac{1-e^{-\beta_{H}\omega}}{e^{-\beta_{H}\omega}-e^{-(\beta_{H}+\beta_{C})\omega}},$ (31)
respectively. Therefore, one must follow the protocol in Box 1 to achieve the best work production and efficiency of the stroke-based discrete heat engine with the hot and cold heat baths at inverse temperatures $\beta_{H}$ and $\beta_{C}$, respectively.
{mdframed}
[roundcorner=4pt,backgroundcolor=blue!5]
List of myfloats 1 Protocol leading to optimal work production and efficiency
  1. 1.
    Take a two level system in the state
    $\rho=\begin{pmatrix}\frac{1}{1+e^{-(\beta_{H}+\beta_{C})\omega}}&0\\ 0&\frac{e^{-(\beta_{H}+\beta_{C})\omega}}{1+e^{-(\beta_{H}+\beta_{C})\omega}}\end{pmatrix}$ (32)
  2. 2.
    Apply a thermal operation that corresponds to the Gibbs stochastic matrix
    $T^{{\cal E}_{H}^{*}}=\begin{pmatrix}1-e^{-\beta_{H}\omega}&1\\ e^{-\beta_{H}\omega}&0\end{pmatrix}$ (33)
    wrt. the hot heat bath.
  3. 3.
    Apply the permutation
    $X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ (34)
    on the system to extract work from the system and transfer it to the battery.
  4. 4.
    Apply a thermal operation that corresponds to the Gibbs stochastic matrix
    $T^{{\cal E}_{C}^{*}}=\begin{pmatrix}1-e^{-\beta_{C}\omega}&1\\ e^{-\beta_{C}\omega}&0\end{pmatrix}$ (35)
    wrt. the cold heat bath which closes the cycle.
From Eq. (30), one can also recover the efficiency of the open cycle engine that we have described in subsection II.3. Since, in the open cycle engine, the working body starts with a thermal state at the temperature of the cold bath $\beta_{C}$, closing the cycle during the final stroke reduces to re-thermalization at the inverse temperature $\beta_{C}$. Therefore, by substituting
$\lambda_{H}^{\text{max}}=1\,\,\textrm{and}\,\,\lambda_{C}^{\text{max}}=\frac{1}{1+e^{-\beta_{C}\omega}}$ (36)
in Eqs. (29) and (30), we obtain [20]:
$W_{\text{max}}=\omega\Big{(}\frac{2e^{-\beta_{H}\omega}}{1+e^{-\beta_{C}\omega}}-1\Big{)},\,\,\mbox{and}$ (37)
$\eta_{\text{max}}=1-\frac{1-e^{-\beta_{H}\omega}}{e^{-\beta_{H}\omega}-e^{-\beta_{C}\omega}}.$ (38)
Next, we shall focus on characterizing the efficiency of the three-stroke engine when we do not have access to a full set of thermal operations. Namely, we focus on two such scenarios: first, we consider the case where we study the optimal performance of the engine when we restrict ourselves to thermal operations that can be realized via the Jaynes-Cummings interaction, and second, when we restrict ourselves to thermal operations that can be performed when one has access only to a finite-dimensional heat bath that can be modeled by a truncated harmonic oscillator.
Figure 2: Plot of efficiency vs $\beta_{C}/\beta_{H}$ with $\omega=0.2/\beta_{H}$. The solid, dashed, dotdashed, and dotted curves represent optimal efficiency, efficiencies for finite heat bath with $d=15$, $d=10$, and $d=5$, respectively.
Figure 3: Plot of $\beta_{H}$ork vs $\beta_{C}/\beta_{H}$ with $\omega=0.2/\beta_{H}$. The solid, dashed, dotdashed, and dotted curves represent optimal work production, work productions for finite heat bath with $d=15$, $d=10$, and $d=5$, respectively.
III.2.1 Optimal performance of the engine for finite sized heat baths
In this section, we explore the engine’s optimal efficiency and work output when we only have access to finite-dimensional heat baths that can be modeled using a truncated harmonic oscillator. In this scenario, one cannot perform any arbitrary thermal operation in the first and third strokes, which implies $\lambda^{\text{max}}<1$ (See Eq. (22)). We need to figure out the value of $\lambda^{\text{max}}$ when we have access only to the finite-dimensional heat baths modeled by a truncated harmonic oscillator. This is addressed in the following theorem.
Theorem 3.
Consider a thermal operation $\mathcal{E}$ on a two-level system given in the state $\rho$
$\mathcal{E}(\rho)=\text{Tr}_{E}\Big{(}U(\rho\otimes\gamma_{E})U^{\dagger}\Big{)},$ (39)
where $\rho$ is a energy incoherent state, i.e., $[\rho,H_{S}]=0$, and $[U,H_{S}+H_{E}]=0$. Let us model the Hamiltonian of the system as
$H_{S}=\omega|1\rangle\langle 1|,$ (40)
which is resonant with the bath Hamiltonian modeled by a truncated harmonic oscillator, i.e.,
$H_{E}=\omega\sum_{n=0}^{d}n|n\rangle\langle n|,$ (41)
where $d+1$ is the dimension of the heat bath. Then, the maximum value of $\lambda$ in Eq. (21) is given by
$\lambda^{\textnormal{max}}=\frac{1-e^{-\beta\omega d}}{1-e^{-\beta\omega(d+1)}}.$ (42)
The proof of this theorem can be found in Appendix D. One can clearly see from the above Eq. (42) that when $d\rightarrow\infty$, then $\lambda^{\text{max}}\rightarrow 1$, which means any thermal operation on a two-level system can be realized with the bath Hamiltonian modeled by a harmonic oscillator. Another interesting limit is when $d\rightarrow 1$, which implies the heat bath is modeled by a two-level system (See Eq. (41)). In that case, we have
$\lambda^{\text{max}}=\frac{1}{1+e^{-\beta\omega}},$ (43)
and substituting Eq. (43) into Eq. (20), we obtain
$\mathcal{T}(\rho)=\text{Conv}\Big{\{}\rho,\frac{1}{1+e^{-\beta\omega}}\begin{pmatrix}1&0\\ 0&e^{-\beta\omega}\end{pmatrix}\Big{\}}.$ (44)
This suggests that when a two-level system functions as a heat bath, it restricts thermal operations between an identity map and thermalization. As a result, when we model a two-level system resonant with the working body as a heat bath, the three stroke engine cannot extract work since the heat stroke fails to induce non-passivity in the working body.
Employing the result from theorem 5 of Appendix D, the values of $\lambda^{\textrm{max}}$ for the inverse temperatures $\beta_{C}$ and $\beta_{H}$ reduce to
$\lambda^{\text{max}}_{C,d}:=\frac{1-e^{-\beta_{C}\omega d}}{1-e^{-\beta_{C}\omega(d+1)}},\,\,\mbox{and}\,\,\lambda^{\text{max}}_{H,d}:=\frac{1-e^{-\beta_{H}\omega d}}{1-e^{-\beta_{H}\omega(d+1)}},$ (45)
respectively. Substituting these values in Eq. (30), we obtain
$\eta_{d}=1-\frac{(1-e^{-d\beta_{C}\omega})(1-e^{-\beta_{H}\omega})(1-e^{-(\beta_{C}+d\beta_{H})\omega})}{(1-e^{-\beta_{C}\omega})(e^{-\beta_{H}\omega}-e^{-d\beta_{C}\omega})(1-e^{-d\beta_{H}\omega})}.$ (46)
Similarly, in this scenario, one can obtain an expression for optimal work production ($W_{d}$) by substituting the values of $\lambda$s given in Eq. (45) into Eq. (29). We plot the work production and efficiency in Figs. 2 and 3 for $\omega=0.2/\beta_{H}$ and some $d$ values, and compared them with $W_{\max}$ and $\eta_{\max}$, respectively.
III.2.2 Optimal performance of the engine modelled with Jaynes-Cummings interaction
In this section, we explore the Jaynes-Cummings model and determine the optimal work and efficiency for the stroke-based discrete engine we considered. Using Eq. (22), we obtain the maximum achievable values of $\lambda$s for hot and cold heat baths as follows[65]:
$\lambda^{\textrm{max}}_{H,\text{JC}}=\begin{cases}\frac{1}{16}\left(8e^{-\beta_{H}\omega}-e^{-2\beta_{H}\omega}+e^{3\beta_{H}\omega}+8\right),&\text{for }0\leq\beta_{H}\omega\leq\frac{\log 4}{3},\\ e^{-4\beta_{H}\omega}-e^{-3\beta_{H}\omega}+1,&\text{for }\beta_{H}\omega\geq\frac{\log 4}{3},\end{cases}$ (47)
and
$\lambda^{\textrm{max}}_{C,\text{JC}}=\begin{cases}\frac{1}{16}\left(8e^{-\beta_{C}\omega}-e^{-2\beta_{C}\omega}+e^{3\beta_{C}\omega}+8\right),&\text{for }0\leq\beta_{C}\omega\leq\frac{\log 4}{3},\\ e^{-4\beta_{C}\omega}-e^{-3\beta_{C}\omega}+1,&\text{for }\beta_{C}\omega\geq\frac{\log 4}{3}.\end{cases}$ (48)
Figure 4: Plot of efficiencies vs $\beta_{C}/\beta_{H}$ with $\omega=0.2/\beta_{H}$. The solid, dashed, dotdashed, and dotted curves represent Carnot efficiency, optimal efficiency, efficiency for finite heat bath with $d=5$, and efficiency for Jaynes-Cummings model, respectively.
Figure 5: Plot of $\beta_{H}$ork vs Efficiency. The solid, dotdashed, dashed, and dotted curves represent optimal work, work for finite heat bath with $d=10$, $d=5$, and work for Jaynes-Cummings, respectively.
Now, substituting $\lambda^{\textrm{max}}_{H,\text{JC}}$ and $\lambda^{\textrm{max}}_{C,\text{JC}}$ from Eqs. (47) and (48) into Eqs. (29) and (30), we can obtain the optimal values of work and efficiency for the engine. The resulting expressions are cumbersome, hence we opted not to put them here; rather, we compare the efficiency obtained in this case with the other cases as shown in Fig. 4. Furthermore, we also make a comparison of the tradeoff between work and efficiency in Fig. 5 for various scenarios we have discussed so far.
IV Conclusion
In this paper, we studied a stroke-based engine in the microscopic regime that requires three strokes to complete the thermodynamic cycles. In particular, we considered the three-stroke engine that involves a two-level system as a working body that steers the energy flow from a hot heat bath to a cold heat bath. The transformation of the working body in the presence of both heat baths is described by thermal operations. We figured out the optimal protocol that leads to the optimal production of work as well as the optimal efficiency when one can perform a restricted class of thermal operations on the working body in the presence of hot and cold heat baths. We analyzed the optimal efficiency and work production by the engine for two such restricted classes: first, when one has access to finite-sized heat baths, and second, when the thermal operations involved in the first and third strokes are realized via the Jaynes-Cummings interaction. Furthermore, we showed that the obtained expressions of optimal work production and efficiency reduce to the expressions of optimal efficiency and work production given in Theorem 3 of Ref. [66] when we relax the restrictions, hence allowing any arbitrary thermal operation on the working body in the presence of hot and cold baths.
Our research paves the way for numerous exciting research directions. One could explore how the optimal efficiency and maximum work produced by the engine change with the dimension of the working body. Exploring how the engine’s optimal efficiency and maximum work production vary with the dimensions of the working body presents an intriguing challenge. The complexity of this problem lies in the intricate structure of the thermal cone, which becomes increasingly convoluted as the dimension of the working body grows. On top of that, the complexity inherent in the Hamiltonian structure of the working body further adds to the intricacy of the problem.
Investigating the significance of coherence in the operation of three-stroke engines presents a captivating avenue for research. In Ref. [66], it was shown that coherence in the state of the working body, modeled as a two-level system, fails to yield performance enhancements in three-stroke engines. This raises the question: is it possible to generalize this result? Specifically, it will be interesting to prove or disprove that the coherence present in the state of the working body is not helpful to enhance the performance of a three-stroke engine. However, characterizing the transformation of coherence in a state under thermal operations remains a formidable challenge, particularly beyond two dimensions. Nonetheless, related results suggest that the dephasing of the state in the local energy eigenbasis is compatible with thermal operations across all dimensions. This result might be helpful in this endeavor.
This work centers on stroke-based heat engines, where work extraction and heat exchange happen in discrete, distinct time intervals. We anticipate a correspondence between the three-stroke discrete heat engine and the continuous heat engine. Notably, analyzing the continuous heat engine enables us to compute the engine’s power or rate of work extraction with time, while the proposed three-stroke heat engine yields the average work per cycle. Upon rigorously bridging the gap between these stroke-based discrete and continuous heat engines, an interesting question emerges: what is the minimum time required to complete a thermodynamic cycle? We envision that elucidating the connection between the three-stroke heat engine and the continuous heat engine could offer valuable insights into realizing the three-stroke heat engines experimentally. In a related study, as proposed in Ref. [67], an experimental implementation of an autonomous refrigerator within the framework of cavity quantum electrodynamics was suggested. Moreover, recent work outlined in Ref. [33] explores the construction of an autonomous refrigerator using superconducting circuits. Therefore, we believe that designing the three-stroke heat engine is attainable within the frameworks of cavity QED and superconducting circuits.
Acknowledgement
TB acknowledges Paweł Mazurek for insightful discussions and comments. The part of the work done at Los Alamos National Laboratory (LANL) was carried out under the auspices of the US DOE and NNSA under contract No. DEAC52-06NA25396. TB also acknowledges support by the DOE Office of Science, Office of Advanced Scientific Computing Research, Accelerated Research for Quantum Computing program, Fundamental Algorithmic Research for Quantum Computing (FAR-QC) project. TB also acknowledge support from the Foundation for Polish Science through IRAP project co-financed by EU within the Smart Growth Operational Programme (contract no.2018/MAB/5). CD acknowledges support from the German Federal Ministry of Education and Research (BMBF) within the funding program “quantum technologies – from basic research to market” in the joint project QSolid (grant number 13N16163).
Appendix A Upper bounding the efficiency of the three stroke engine using Carnot efficiency
Lemma 4.
For any discrete stroke-based heat engine, if the work production $W$ is positive, then the amount of heat flowing from the hot heat bath $Q_{H}$ is also positive, i.e.,
$W>0\Rightarrow Q_{H}>0.$ (49)
Furthermore, the efficiency of the discrete stroke-based engine will be upper bounded by Carnot efficiency, i.e.,
$\eta=\frac{W}{Q_{H}}\leq 1-\frac{\beta_{H}}{\beta_{C}}=\eta_{\textnormal{Carnot}}.$ (50)
Proof.
From Eq. (7) of section II.2, we know that
$W=\text{Tr}\Big{(}H_{S}\big{(}\mathcal{E}_{H}(\rho_{S})-V(\mathcal{E}_{H}(\rho_{S}))V^{\dagger}\big{)}\Big{)}.$ (51)
and we rewrite the transfer from the hot and cold baths as Eq. (5) and cold heat Eq. (13)
$Q_{H}=\text{Tr}\Big{(}\big{(}H_{S}\big{(}\mathcal{E}_{H}(\rho_{S})-\rho_{S}\big{)}\Big{)}\,\,\mbox{and}$ (52)
$Q_{C}=\text{Tr}\Big{(}H_{C}\big{(}{\cal E}_{C}(V{\cal E}_{H}(\rho_{S})V^{\dagger})-V{\cal E}_{H}(\rho_{S})V^{\dagger}\big{)}\Big{)}$
$\,\,\,\,\,\,\,\,\,\,=\text{Tr}\Big{(}H_{C}\big{(}\rho_{S}-V{\cal E}_{H}(\rho_{S})V^{\dagger}\big{)}\Big{)}$ (53)
respectively, where the second equality in Eq. (53) follows from the cyclicity condition, i.e.,
${\cal E}_{C}(V{\cal E}_{H}(\rho_{S})V^{\dagger})=\rho_{S}$ (54)
Therefore, we have
$Q_{H}+Q_{C}>0.$ (55)
Next, using the monotonicity of free energy under thermal operations, we can write the following:
$F(\rho_{S})-F(\mathcal{E}_{H}(\rho_{S}))\geq 0,$ (56)
where $F(\cdot)=\text{Tr}(H_{S}\cdot)-S(\cdot)/\beta_{H}$ where $S(\cdot)$ denotes the von Neumann entropy. Thus, one can simplify Eq. (56) as
$\text{Tr}(H_{S}\rho_{S})-\text{Tr}\big{(}H_{S}\mathcal{E}_{H}(\rho_{S})\big{)}-\frac{1}{\beta_{H}}\big{(}S(\rho_{S})-S(\mathcal{E}_{H}(\rho_{S}))\big{)}\geq 0$ (57)
$\Rightarrow$ $-\frac{1}{\beta_{H}}\big{(}S(\rho_{S})-S(\mathcal{E}_{H}(\rho_{S}))\big{)}\geq\text{Tr}(H_{S}\rho_{S})-\text{Tr}\big{(}H_{S}\mathcal{E}_{H}(\rho_{S})\big{)}=Q_{H}$
$\Rightarrow$ $-\Big{(}S(\rho_{S})-S(\mathcal{E}_{H}(\rho_{S}))\Big{)}\geq\beta_{H}Q_{H}.$
Similarly, for the cold heat bath, we find
$-\Big{(}S(V\mathcal{E}_{H}(\rho_{S})V^{\dagger})-S\big{(}\mathcal{E}_{C}(V\mathcal{E}_{H}(\rho_{S})V^{\dagger})\big{)}\Big{)}\geq\beta_{C}Q_{C}$ (58)
$\Rightarrow$ $-\Big{(}S(\mathcal{E}_{H}(\rho_{S}))-S(\rho_{S})\Big{)}\geq\beta_{C}Q_{C},$
where we have use the fact $S(V\mathcal{E}_{H}(\rho_{S})V^{\dagger})=S(\mathcal{E}_{H}(\rho_{S}))$ as von-Neumann entropy is invariant under unitary operation, and cyclicity condition given in Eq. (54) to write Eq. (58). Now, adding Eqs. (57) and (58), we obtain $\beta_{H}Q_{H}+\beta_{C}Q_{C}\leq 0$, that gives
$-\frac{\beta_{H}}{\beta_{C}}Q_{H}\geq Q_{C}.$ (59)
Substituting this upper bound of $Q_{C}$ given in Eq. (59) into Eq. (55) and noting the fact that $\beta_{H}<\beta_{C}$, we arrive at
$Q_{H}\Big{(}1-\frac{\beta_{H}}{\beta_{C}}\Big{)}>0\quad\Rightarrow\;Q_{H}>0,$ (60)
which completes the proof of the first part. As a consequence, employing Eq. (59) we obtain
$\eta=\frac{W}{Q_{H}}=\frac{Q_{H}+Q_{C}}{Q_{H}}=1+\frac{Q_{C}}{Q_{H}}\leq 1-\frac{\beta_{H}}{\beta_{C}},$ (61)
which completes the proof for the second part. ∎
Appendix B Thermal operation and $\beta$-ordering
We have introduced a thermal operation on a given state $\rho_{S}$ of dimension $d$ as
${\cal E}(\rho_{S})=\text{Tr}_{E}\left({U\left(\rho_{S}\otimes\gamma_{E}\right)U^{\dagger}}\right),$ (62)
where $U$ is a joint unitary acting on the system and a heat bath that commutes with the total Hamiltonian of the system and bath, i.e., $[U,H_{S}+H_{E}]=0$, and $\gamma_{E}$ is a thermal state of the bath with very large volume at some fixed inverse temperature $\beta$. If $\rho_{S}$ is an energy-incoherent state, i.e., $[\rho_{S},H_{S}]=0$, one can prove [26]
$\rho_{S}\xrightarrow{{\cal E}}{\cal E}(\rho_{S}):=\sigma_{S}\quad\Longrightarrow\quad T^{{\cal E}}(\bm{p})=\bm{q},\quad T^{{\cal E}}(\bm{\gamma})=\bm{\gamma},$ (63)
where $\bm{p}$ , $\bm{q}$ and $\bm{\gamma}$ are the population vectors in $d$-dimension constructed from $\rho_{S}$ , ${\cal E}(\rho_{S})$ and the Gibbs state of the system $\gamma$, respectively. Recall that we defined a population vector for any state of the system $\rho$ as a probability vector with entries $p_{i}=\langle{E_{i}}|\rho|{E_{i}}\rangle$, where $\{|{E_{i}}\rangle\}$s correspond to the eigenbasis of the Hamiltonian of the system.
Employing the concepts of relative majorization, one can verify that the existence of such a Gibbs-stochastic matrix $T^{\mathcal{E}}$ is equivalent to the thermomajorization condition $\bm{p}\succ_{\gamma}\bm{q}$. For a detailed review of thermomajorization, we refer to Ref. [55, 62, 63]. Here, we will outline the ideas of the thermonajorization condition using the concept of the thermomajorization curve, which is necessary to analyze the engine’s performance.
[…]
$\frac{p_{\pi(1)}}{\gamma_{\pi(1)}}\geq\frac{p_{\pi(2)}}{\gamma_{\pi(2)}}\ldots\geq\frac{p_{\pi(d)}}{\gamma_{\pi(d)}},$ (64)
and $Z_{S}$ is the partition function of the system obtained from the Hamiltonian $H_{S}$. The curve $L_{\bm{\gamma}}(\bm{p})$ is referred to as the thermomajorization curve of $\bm{p}$ with respect to $\bm{\gamma}$, and the ordering obtained by the action of permutation $\pi$ on the set of indices $(12\ldots d)$, i.e.,
$\pi(12\ldots d),$ (65)
is referred to as $\beta$-order of $\bm{p}$ with respect to $\bm{\gamma}$. We say $\bm{p}$ thermomajorizes $\bm{q}$ and denote it as $\bm{p}\succ_{\gamma}\bm{q}$ if and only if the thermomajorization curve $L_{\bm{\gamma}}(\bm{p})$ does not lie below the thermomajorization curve $L_{\bm{\gamma}}(\bm{q})$. Therefore, the existence of a thermal operation that transforms $\rho$ to $\sigma$ is equivalent to the existence of a Gibbs-stochastic matrix that transforms the population vector $\bm{p}$ associated with $\rho$, to the population vector $\bm{q}$ associated with $\sigma$, and this further reduces to the condition $\bm{p}\succ_{\gamma}\bm{q}$. In the limit $\beta\rightarrow 0$, $\bm{p}\succ_{\gamma}\bm{q}$ boils down to the well-known majorization condition $\bm{p}\succ\bm{q}$.
Appendix C Proof of theorem 2
We begin by re-writing the population vector associated with the initial state $\rho_{S}$ as
$\bm{p}=\begin{pmatrix}p&1-p\end{pmatrix}^{T}.$ (66)
Since we are working with a two-level system as a working body, it can have only two possible $\beta$-order wrt. inverse temperature $\beta_{H}$:
$(12)\quad\text{when}\quad p\geq\frac{1-p}{e^{-\beta_{H}\omega}},$ (67)
$(21)\quad\text{when}\quad p\leq\frac{1-p}{e^{-\beta_{H}\omega}}.$ (68)
If the population vector of the initial state of the working body has $\beta$-order $(21)$, then it has the highest average energy among all states that can be achieved from the initial state $\rho_{S}$ via thermal operation at inverse temperature $\beta$ (See Lemma 16 of Ref. [20]). Therefore, if the population vector of the initial state of the working body has $\beta$-order $(21)$, then the change in the average energy under any thermal operation ${\cal E}_{\beta}(\cdot)$ at inverse temperature $\beta$ will either be zero or negative
$\text{Tr}(H_{S}{\cal E}_{H}(\rho_{S}))-\text{Tr}(H_{S}\rho_{S})\leq 0.$ (69)
Recall that $\text{Tr}(H_{S}{\cal E}_{H}(\rho_{S}))-\text{Tr}(H_{S}\rho_{S})$ gives an exchanged amount of heat from the heat bath at temperature $\beta_{H}$ (see Eq. (5)), which will be negative if the $\beta$-order of the initial state $\rho_{S}$ is negative. Employing the lemma 4 further implies that the work done $W$ must be negative, which implies the engine cannot function if the initial state of the working body $\rho_{S}$ has a $\beta$-order $(21)$.
Therefore, in order to extract work, the initial state of the two-level working body should have $\beta$-order $(12)$ wrt. the hot heat bath, which allows us to write from Eq. (67) the following:
$p\geq\frac{1-p}{e^{-\beta_{H}\omega}}\;\;\Rightarrow\;\;e^{-\beta_{H}\omega}\geq\frac{1-p}{p}.$ (70)
Then, we can re-parameterize the state of the two-level working body $\bm{p}$ using the effective temperature $\beta_{C}^{v}$ defined by the following relation [12]:
$\frac{1-p}{p}=e^{-\beta_{C}^{v}\omega}.$ (71)
Using this parametrization, we can compute $p=\frac{1}{1+e^{-\beta_{C}^{v}\omega}}$. Furthermore, using Eqs. (70) and (71), we can easily see $\beta_{C}^{v}\geq\beta_{H}$.
Here, we would like to make a crucial remark. We can imagine any three-stroke engine with a two-level working body and population vector $\bm{p}$ associated with two thermal baths at inverse temperatures $\beta_{H}$ and $\beta_{C}$ as an open-cycle engine that is associated with two thermal baths at inverse temperatures $\beta_{H}$ and $\beta^{v}_{C}$ (Recall that an open cycle engine is a three-stroke engine where the final stroke that aims to close the thermodynamic cycle is given by thermalization, as described in Sec. II.3. We illustrate this idea in Fig. 1.) In other words, any stroke-based discrete heat engine with a two-level working body is essentially an open-cycle heat engine wrt. a new effective inverse temperature of the cold heat bath.
Now, theorem 6 of Ref. [20] states that the optimal work output and efficiency for an open cycle engine are obtained at the extremal points of the thermal cone $\mathcal{T}_{H}(\bm{p})$, where the work extracting unitary in the work stroke is provided by
$X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},$ (72)
where ${\cal T}_{H}(\bm{p})$ is the thermal cone at the inverse temperature $\beta_{H}$ which justifies the subscript (Recall that thermal cone is defined as set of all states that can be achieved from $\rho_{S}$ via thermal operation at inverse temperature $\beta$, as described in Sec. II.4). As a result of the correspondence between the three-stroke engine and the open-cycle engine, we do not need to optimize the three-stroke engine’s performance over the heat and work strokes.
Now, given a fixed $\lambda^{\text{max}}_{H}\in[0,1]$, the extremal points of $\mathcal{T}_{H}(\bm{p})$ are given by
$\Big{\{}\bm{p},\lambda_{H}^{\text{max}}\begin{pmatrix}1-pe^{-\beta_{H}\omega}\\ pe^{-\beta_{H}\omega}\end{pmatrix}+(1-\lambda_{H}^{\text{max}})\bm{p}\Big{\}},$ (73)
where the second extremal point can only lead to non-trivial work production and efficiency. Observe that, when $\lambda_{H}^{\text{max}}=1$, we obtain the full thermal cone as given in Ref. [26].
Therefore, optimal work production and efficiency have the following forms
$W^{*}(p)$ $=$ $\omega\left(1-2\lambda_{H}^{\text{max}}+2p(\lambda_{H}^{\text{max}}e^{-\beta_{H}\omega}+\lambda_{H}^{\text{max}}-1)\right),$ (74)
$\eta^{*}(p)$ $=$ $1-\frac{2p-1-\lambda_{H}^{\text{max}}\left(pe^{-\beta_{H}\omega}+p-1\right)}{\lambda_{H}^{\text{max}}(pe^{-\beta_{H}\omega}+p-1)}.$ (75)
From Eq. (74), it is straightforward to see that $W^{*}$ is an increasing function of $p$. On the other hand, by differentiating $\eta^{*}$ given in Eq. (75) wrt. $p$, one can also see that $\eta^{*}(p)$ is an increasing function of $p$:
$\frac{\partial\eta^{*}(p)}{\partial p}=\frac{1-e^{-\beta_{H}\omega}}{\lambda^{\text{max}}_{H}\big{(}pe^{-\beta_{H}\omega}+p-1\big{)}^{2}}>0.$ (76)
Therefore, here our interest is to find out the highest possible value of $p$ that is compatible with the cyclicity condition. In order to do so, we proceed as follows.
As the optimal efficiency and work produced by the engine achieved at the extremal point of the set ${\cal T}_{H}(\bm{p})$, thus from Eq. (73), one can see that the population vector of the $\bm{p}$ associated with the initial state transformed to the population vector
$\lambda_{H}^{\text{max}}\begin{pmatrix}1-pe^{-\beta_{H}\omega}\\ pe^{-\beta_{H}\omega}\end{pmatrix}+(1-\lambda_{H}^{\text{max}})\bm{p}=\lambda_{H}^{\text{max}}\begin{pmatrix}1-pe^{-\beta_{H}\omega}\\ pe^{-\beta_{H}\omega}\end{pmatrix}+(1-\lambda_{H}^{\text{max}})\begin{pmatrix}p\\ 1-p\end{pmatrix},$ (77)
during the heat stroke. After the work stroke which is governed by the unitary $X$ in Eq. (72), the resulting population vector from the heat stroke given in Eq. (77) transformed to
$\lambda_{H}^{\text{max}}\begin{pmatrix}pe^{-\beta_{H}\omega}\\ 1-pe^{-\beta_{H}\omega}\end{pmatrix}+(1-\lambda_{H}^{\text{max}})\begin{pmatrix}1-p\\ p\end{pmatrix}=\begin{pmatrix}\lambda_{H}^{\text{max}}pe^{-\beta_{H}\omega}+(1-\lambda^{\text{max}}_{H})p\\ \lambda_{H}^{\text{max}}\left(1-pe^{-\beta_{H}\omega}\right)+(1-\lambda^{\text{max}}_{H})(1-p)\end{pmatrix}:=\begin{pmatrix}z\\ 1-z\end{pmatrix}=\bm{z},$ (78)
where $z=\lambda_{H}^{\text{max}}pe^{-\beta_{H}\omega}+(1-\lambda^{\text{max}}_{H})p$. Now the final stroke which is a thermal operation wrt. a cold bath, should transform the population vector $\bm{z}$ to the initial state which results in closing the cycle. Therefore,
$\lambda_{C}\begin{pmatrix}1-ze^{-\beta_{C}\omega}\\ ze^{-\beta_{C}\omega}\end{pmatrix}+(1-\lambda_{C})\begin{pmatrix}z\\ 1-z\end{pmatrix}=\begin{pmatrix}p\\ 1-p\end{pmatrix}.$ (79)
By substituting $z$ from Eq. (78) in Eq. (79) and solving the resulting equation for $p$, we obtain
$p=\frac{1-\lambda^{\text{max}}_{H}(1-\lambda_{C})-\lambda_{C}(1-\lambda^{\text{max}}_{H})e^{-\beta_{C}\omega}}{2-\lambda_{C}(1-\lambda_{H}^{\text{max}})-\lambda_{H}^{\text{max}}-\lambda_{C}(1-\lambda_{H}^{\text{max}})e^{-\beta_{C}\omega}-\lambda_{H}^{\text{max}}(1-\lambda_{C})e^{-\beta_{H}\omega}+\lambda_{H}^{\text{max}}\lambda_{C}e^{-(\beta_{C}+\beta_{H})\omega}}.$ (80)
We can see that $p$ is an increasing function of $\lambda_{C}$ by differentiating $p$ wrt. $\lambda_{C}$ as
$\frac{\partial p}{\partial\lambda_{C}}=\frac{e^{\omega(\beta_{C}+\beta_{H})}\Big{(}e^{\omega(\beta_{C}+\beta_{H})}-\big{(}\lambda^{\text{max}}_{H}e^{\beta_{C}\omega}+(1-\lambda^{\text{max}}_{H})e^{\beta_{H}\omega}\big{)}\Big{)}}{\bigg{(}\Big{(}2-\lambda^{\text{max}}_{H}-(1-\lambda_{H}^{\text{max}})\lambda_{C}\Big{)}e^{\omega(\beta_{C}+\beta_{H})}-(1-\lambda_{C})\lambda_{H}^{\text{max}}e^{\beta_{C}\omega}-(1-\lambda_{H}^{\text{max}})\lambda_{C}e^{\beta_{H}\omega}+\lambda_{C}\lambda_{H}^{\text{max}}\bigg{)}^{2}}>0.$ (81)
The inequality in the above Eq. (81) follows because $e^{\beta_{C}\omega}>1$ and $e^{\beta_{H}\omega}>1$, therefore, the second term in the numerator appeared on RHS of Eq. (81) is positive, i.e.,
$e^{\omega(\beta_{C}+\beta_{H})}-\big{(}\lambda^{\text{max}}_{H}e^{\beta_{C}\omega}+(1-\lambda^{\text{max}}_{H})e^{\beta_{H}\omega}\big{)}>0,$ (82)
which implies that the numerator is always positive, whereas the denominator is always positive because of the square. Therefore, the largest possible value of $p$ compatible with the cyclicity condition is obtained when $\lambda_{C}=\lambda_{C}^{\text{max}}$, i.e.,
$p=\frac{1-\lambda^{\text{max}}_{H}(1-\lambda_{C}^{\text{max}})-\lambda_{C}^{\text{max}}(1-\lambda^{\text{max}}_{H})e^{-\beta_{C}\omega}}{2-\lambda_{C}^{\text{max}}(1-\lambda_{H}^{\text{max}})-\lambda_{H}^{\text{max}}-\lambda_{C}^{\text{max}}(1-\lambda_{H}^{\text{max}})e^{-\beta_{C}\omega}-\lambda_{H}^{\text{max}}(1-\lambda_{C}^{\text{max}})e^{-\beta_{H}\omega}+\lambda_{H}^{\text{max}}\lambda_{C}^{\text{max}}e^{-(\beta_{C}+\beta_{H})\omega}}.$ (83)
Substituting the value of $p$ from Eq. (83) in Eqs. (74) and (75), we get
$W_{\text{max}}$ $=$ $1-2\lambda^{\text{max}}_{H}+\frac{2\Big{(}\lambda_{H}e^{-\beta_{H}\omega}-(1-\lambda_{H}^{\text{max}})\Big{)}\Big{(}1-(1-\lambda_{C}^{\text{max}})\lambda_{H}^{\text{max}}-\lambda_{C}^{\text{max}}e^{-\beta_{C}\omega}(1-\lambda_{H}^{\text{max}})\Big{)}}{2-\lambda_{C}^{\text{max}}(1-\lambda_{H}^{\text{max}})-\lambda_{H}^{\text{max}}-\lambda_{C}^{\text{max}}e^{-\beta_{C}\omega}(1-\lambda_{H}^{\text{max}})-\lambda_{H}^{\text{max}}e^{-\beta_{H}\omega}(1-\lambda_{C}^{\text{max}})+\lambda_{C}^{\text{max}}\lambda_{H}^{\text{max}}e^{-(\beta_{C}+\beta_{H})\omega}},$
$\eta_{\text{max}}$ $=$ $1-\frac{\lambda_{C}^{\text{max}}\Big{(}1-\lambda_{H}^{\text{max}}e^{-\beta_{H}\omega}-(1-\lambda_{H}^{\text{max}})e^{-\beta_{C}\omega}\Big{)}}{\lambda^{\text{max}}_{H}\Big{(}e^{-\beta_{H}\omega}-(1-\lambda^{\text{max}}_{C})-\lambda_{C}^{\text{max}}e^{-(\beta_{H}+\beta_{C})\omega)}\Big{)}}.$ (84)
This completes the proof of theorem 2.
Appendix D Range of thermal operation with finite sized baths
In this section, we describe the range of thermal operations for a qubit system with a finite size heat bath. The thermal operations on a qubit system can be expressed as a convex mixture of two extremal thermal processes $\lambda T^{{\cal E}}+(1-\lambda)\mathbbm{1}$, where $0\leq\lambda\leq 1$ and
$\mathcal{E}=\begin{pmatrix}1-e^{-\beta\omega}&1\\ e^{-\beta\omega}&0\end{pmatrix}.$ (85)
However, when the bath is of finite dimension the full range of $\lambda$ is not allowed. Let’s consider the dimension of the heat bath is $d_{E}=d+1$. Tha Hamiltonians of the system and heat bath are
$H_{S}=\omega|{1}\rangle\!\langle{1}|\,\,\mbox{and}\,\,H_{E}=\omega\sum_{n=0}^{d}n|{n}\rangle\!\langle{n}|$ (86)
respectively. Note that we assume that the ground state energies of the system and heat bath are zero, and the energy levels are eqispaced. With this we can state the following theorem:
Theorem 5.
The allowed range of thermal operation for a qubit system when the heat bath is finite, is $\lambda\in\{0,\lambda^{\max}\}$, where
$\lambda^{\max}=\frac{1-e^{-\beta\omega d}}{1-e^{-\beta\omega(d+1)}}.$ (87)
Proof.
As $[U,H_{S}\otimes\mathbbm{1}_{E}+\mathbbm{1}_{S}\otimes H_{E}]=0$, then for a thermal operation with such a heat bath, $U$ has the following $2\times 2$ block-diagonal structure in the energy eigenbasis:
$U=\mathbbm{1}_{\{|{0}\rangle_{S}|{0}\rangle_{E}\}}\bigoplus_{i=1}^{d}V_{i_{\{|{0}\rangle_{S}|{i}\rangle_{E},|{1}\rangle_{S}|{i-1}\rangle_{E}\}}}\oplus\mathbbm{1}_{\{|{1}\rangle_{S}|{d}\rangle_{E}\}},$ (88)
where $V_{i}$’s are $2\times 2$ unitary matrix in the sub-space $\{|{0}\rangle_{S}|{i}\rangle_{E},|{1}\rangle_{S}|{i-1}\rangle_{E}\}$ and $\mathbbm{1}$’s are basically 1. Further, we take initial system state $\rho_{S}$ to be diagonal in the energy eigenbasis, i.e.,
$\rho_{S}=\begin{pmatrix}p&0\\ 0&1-p\end{pmatrix}.$ (89)
The state of the bath is
$\rho_{E}=\frac{e^{-\beta H_{E}}}{Z_{H_{E}}}=\frac{1}{Z_{H_{E}}}\begin{pmatrix}1&0&\cdots&0\\ 0&e^{-\beta\omega}&\cdots&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&e^{-d\beta\omega}\end{pmatrix},$ (90)
where $Z_{H_{E}}=\mbox{Tr}\left(e^{-\beta H_{E}}\right)$. Hence, the initial state of the system and bath has the following form in the energy eigenbasis (see equation Eq. (88)):
$\rho_{S}\otimes\rho_{E}=\frac{1}{Z_{H_{E}}}=p\bigoplus_{i=1}^{d}\begin{pmatrix}pe^{-i\beta\omega}&0\\ 0&(1-p)e^{-(i-1)\beta\omega}\end{pmatrix}\oplus(1-p)e^{-d\beta\omega}.$ (91)
Note that the energies of the blocks from left to right are $0,\omega,2\omega,\ldots,(d+1)\omega$. As the unitary and the initial state of the system and bath have similar block diagonal structures, $U\left(\rho_{S}\otimes\rho_{E}\right)U^{\dagger}$ also has a similar block diagonal structure. Therefore, it is enough to find out the elements of one block. Consider a block position $j$ with energy $j\omega$, then we have
$\rho_{SE}$ $=$ $V_{j}(\rho_{S}\otimes\rho_{E})V_{j}^{\dagger}=\begin{pmatrix}e^{i\phi_{j}}\cos\theta_{j}&e^{i\alpha_{j}}\sin\theta_{j}\\ -e^{-i\alpha_{j}}\sin\theta_{j}&e^{-i\phi_{j}}\cos\theta_{j}\end{pmatrix}\begin{pmatrix}pe^{-j\beta\omega}&0\\ 0&(1-p)e^{-(j-1)\beta\omega}\end{pmatrix}\begin{pmatrix}e^{i\phi_{j}}\cos\theta_{j}&-e^{i\alpha_{j}}\sin\theta_{j}\\ e^{-i\alpha_{j}}\sin\theta_{j}&e^{i\phi_{j}}\cos\theta_{j}\end{pmatrix}$
$=$ $\begin{pmatrix}pe^{-j\beta\omega}\cos^{2}\theta_{j}+(1-p)e^{-(j-1)\beta\omega}\sin^{2}\theta_{j}&e^{i(\phi_{j}+\alpha_{j})}\left(e^{-(j-1)\beta\omega}(1-p)-pe^{-j\beta\omega}\right)\sin\theta_{j}\cos\theta_{j}\\ e^{-i(\phi_{j}+\alpha_{j})}\left(e^{-(j-1)\beta\omega}(1-p)-pe^{-j\beta\omega}\right)\sin\theta_{j}\cos\theta_{j}&pe^{-j\beta\omega}\sin^{2}\theta_{j}+(1-p)e^{-(j-1)\beta\omega}\cos^{2}\theta_{j}\end{pmatrix},$
where $V_{j}$ is an arbitrary $2\times 2$ special unitary matrix. After tracing out the thermal bath from the final state $\rho_{SE}$, we get the final state of the system as follows
$\rho^{\prime}_{S}=\frac{1}{Z_{H_{E}}}\begin{pmatrix}\mathcal{A}&0\\ 0&Z_{H_{E}}-\mathcal{A}\end{pmatrix},$ (93)
where $\mathcal{A}=p+\sum_{j=1}^{d}\left(pe^{-j\beta\omega}\cos^{2}\theta_{j}+(1-p)e^{-(j-1)\beta\omega}\sin^{2}\theta_{j}\right)$.
On the other hand, under thermal operations as described above theorem 5, one of the diagonal elements of the state $\rho_{S}$ changes to
$p\rightarrow p^{\prime}=1-\left((1-p)(1-\lambda)+e^{-\beta\omega}p\lambda\right).$ (94)
To find the solution of $\lambda$, we need to compare $p^{\prime}$ with $\mathcal{A}/Z_{H_{E}}$, i.e.,
$\lambda=\frac{1}{Z_{H}}\sum_{j=1}^{d}\sin^{2}\theta_{j}e^{-\beta(j-1)\omega}.$ (95)
The maximum value of $\lambda$ is obtained when $\sin^{2}\theta_{j}=1$ $\forall\theta_{j}$ and the value is
$\lambda^{\max}=\sum_{j=1}^{d}\frac{e^{-\beta(j-1)\omega}}{Z_{H}}=\frac{1-e^{-\beta\omega d}}{1-e^{-\beta\omega(d+1)}}.$ (96)
This completes the proof. ∎
Therefore, the range of $\lambda$ is $\{0,\lambda^{\max}\}$. It is easy to check that for $d\rightarrow\infty$, $\lambda^{\max}=1$, and thus all thermal operations are possible. If the heat bath is a qubit, then $\lambda^{\max}=1/\left(1+e^{\beta\omega}\right)$, which implies that the thermal operation corresponds to $\lambda^{\max}$ just thermalizes any input system.